Shift Summary for Owl shift, 07/Jul/99

Peter Shanahan (shanahan@fnal.gov)
Tue, 6 Jul 1999 07:23:30 -0500

Shift Summary for Owl shift, 07/Jul/99

On Shift: Mike Arenton, Peter Shanahan

Hours of beam available at all intensities: 0.5
Hours of beam available at nominal intensity: 0.2
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 6.5

Accounting of lost time: No beam due to Booster Cooling, Booster RF,
and TeV Cryo problems. DAQ was up the entire shift.

========================================================================
Runs ended this shift:
Run 13819 - - LASER_SCAN
Run 13820 - - LASER_SCAN
Run 13821 - - LASER_SCAN
Run 13823 - - COSMIC (Disk only)
Run 13824 - - COSMIC (neither disk nor tape)
Run 13826 - 200 K - BEAM_MUON_4E12

Tapes brought to FCC this shift: QK{S,K,C}x{089-091}, QKB031, QKN007,
QKNL07, QKSM04

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Wait for beam for most of shift. Succeed in keeping DC alcohol
at -5 C.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

- - 00:00 Swap blue ice modules. Left gas shed door open to
help cooldown.
- - 00:30 Load fresh data tapes. Fresh start.
13819 1 01:10 Laser scan with nominal timing
- - 01:30 "Est. TeV startup 5AM." We forget to call Elizabeth.
13820 1 01:56 Laser scan with timing=nominal - 2 ns
13821 1 02:46 Laser scan with timing=nominal + 2 ns
02:50 Alcohol Temperature=-4.8 C
13823 1 03:50 Cosmics with no writing to tape, to excercise DA.
13823 04:25 ONSPLIT requests QKLA06 mount on ktev3. Done.
13823 04:30 "TeV turn on in progress. Beam in 30'."
13824 04:35 Cosmics with no writing to disk or tape.
13826 1 05:27 Start BEAM_MUON run, still no beam
05:40 Swap blue ice modules (3rd time this shift.) T=-4.9C
13826 05:47 "8GeV PS problem, est. 30' for beam."
13826 46 06:30 Beam arrives, but takes ~15' to get about 1E12.
13827 1 07:00 Start E832 data taking.

========================================================================
Follow up Notes on Previous Shift

Blue ice swaps and open gas shed door help keep alcohol temperature at
-4.9C for most of shift.

Tape drive tps2d2 still out of commission on ktev2.

========================================================================
Items to Follow up on Next Shift

Keep swapping blue ice periodically. Temperature today is expected to
be significantly cooler than yesterday, so perhaps there will not be any
problem.